How many grams of pure acid and how many grams of a 40% solution should be mixed together to get a total of 30 grams of a 60% solution?
Answer by mananth(16946)

percent ---------------- quantity
Pure acid 100 ---------------- x grams
Acid 40 ------ 30 - x grams
Mixture 60.00% ---------------- 30

100 x + 40 ( 30 - x ) = 60 * 30

100x+1200-40x= 1800
100 x - 40 x=1800 - 1200
60 x = 600
/ 60
x= 10 grams 100.00% Pure acid

20 grams 40.00% acid